The sum of first three terms of a finite geometric series is -(7/10) and their product is -(1/125). [Hint: Use a/r, a, and ar to represent the first three terms, respectively.] The three numbers are _____, _____, and _____.
Answer by stanbon(75887)

The sum of first three terms of a finite geometric series is -(7/10) and their product is -(1/125). [Hint: Use a/r, a, and ar to represent the first three terms, respectively.] The three numbers are _____, _____, and _____.
------
a/r + a + ar = -7/10
(a/r)*a*(ar) = -1/125
---------------------------
a^3 = -1/125
a = -1/5
------------------------
Solve for "r"::
a/r = (-1/5)/r = -1/(5r)
a = -1/5
a*r = (-1/5)r
------
So, -1/(5r) + (-1/5) + -r/5 = -7/10
Multiply thru by 10r to get:
-2 - 2r -2r^2 = -7r
----
2r^2 -5r + 2 = 0
----
r = 2
-----------------------------------
1st term:: a/r = (-1/5)/(2) = -1/10
2nd term:: a = -1/5 = -2/10
3rd term:: a*r = (-2/10)(2) =-4/10
Note: Sum = -7/10
-------------
Cheers,
Stan H.